# Before & After: The Effect of EU’s 2022 Code of Practice on Disinformation

### v1

@@ 4. Regulation Impact
We set out to explore if the Strengthened Code of Practice on Disinformation has had a meaningful impact on the monetization of misinformation content. We expect that the ecosystem will be improving because multiple signatories have committed to cutting the advertising revenue of disinformation spreaders (codeOfPractice, ). Inspired by previous work (papadogiannakis2023funds, ), we study how the direct business relationships among misinformation websites and ad networks have evolved over time.
Figure 5. Direct business relationships of top ad networks with unpopular misinformation websites over the span of 2 years.
We utilize the dataset of ads.txt and sellers.json files described in Section 3.2, as well as the one collected in (papadogiannakis2023funds, ) for news websites in December 2021. Arguably, publishers of fake news content could easily lie in their ads.txt files (papadogiannakis2023pooling, ; vekaria2023inventory, ). To discover the actual business relationships, we focus on information coming from both the publishers and the ad systems. That is, we cross-reference the claims that publishers make in their ads.txt files with those of trustworthy ad networks in sellers.json files. We argue that since these are conflicting sides of the same coin, the intersection of their claims is the ground truth. For each distinct DIRECT ads.txt entry, we review whether the respective publisher ID is also disclosed by the respective ad network. For instance, if examplefakenews.com has the entry adnetwork.com, 12345, DIRECT in its ads.txt file, we confirm that adnetwork.com also discloses the identifier 12345 in its sellers.json file. Again, we focus only on ads.txt records labeled as DIRECT because they indicate a direct business contract between the misinformation publisher and the advertising system. First, we study the behavior of the top ad networks and how their business relationships with misinformation websites have evolved over a period of 2 years. We focus on the intersection of these datasets and find that there exist 461 misinformation websites with an ads.txt file in both datasets and with at least one business relationship also disclosed by the respective ad network. We plot our findings in Figure 3.
At first glance, we observe that the majority of ad networks still (i.e., at January 2024) work with misinformation websites and that they have not pulled away from them. In fact, 6 out of the 9 most popular ad networks (i.e., rubiconproject.com to openx.com in Figure 3) have only blocked a very small percentage of misinformation websites (i.e., up to 5.4%).
However, one can observe that there exist ad systems that evidently have stopped working with a substantial amount of misinformation websites. Indeed, Google, Lijit and Amazon seem to have dropped 51%, 43% and 37% of misinformation websites that they worked with 2 years ago. Although this drop is significant and seems to be in line with the CoP, a closer look reveals that these websites are insignificant and represent only a tiny portion of the misinformation traffic out there. To further investigate this point, we use the SimilarWeb analytics platform (similarWeb, ) to extract traffic data for misinformation websites. We successfully extract the network traffic of 95% of the misinformation websites for which we have collected their ads.txt files, both before and after the CoP was signed. In Figure 3, we plot the distribution of their network traffic in terms of number of visits on desktop and mobile during January 2024. Please note that the $x$-axis is in logarithmic scale, suggesting that there exist a lot of misinformation websites with low network traffic and a few highly popular that are able to attract millions of monthly visitors. We split these websites on the 80th percentile based on the observation that the top 20% are popular websites, while the rest are unpopular. [^3] The split 80-20 could have been made as 85-15, or 75-25. The point here is not to define an optimal threshold but to create two categories: the “popular” misinformation websites and the “unpopular” ones. The 80th percentile corresponds to a network traffic of approximately 290K monthly visits. Consequently, we find 88 misinformation websites which we label as “popular”, and 351 labeled as “unpopular”.
We perform a similar analysis to the one described in Figure 3 and plot the evolution of business relationships of the top ad networks with popular and unpopular misinformation websites in Figures 5 and 5, respectively. These figures paint a clear picture of the ad ecosystem. That is, even though it seems that some ad networks indeed no longer work with misinformation websites, they do so only for the unpopular ones. In fact, some of these websites have hardly any visitors and their effect on the dissemination of misinformation is limited, deeming them potentially insignificant. When it comes to the big players of misinformation, that are extremely popular and attract up to 100s of millions of visits on a monthly basis, we observe in Figure 5 that ad networks still work with them, even a year and half after the CoP was signed. Few top ad networks no longer have business relationships with a small percentage of the most popular misinformation websites, while we find ad networks such as Amazon and Microsoft (i.e., appnexus.com) that now work with more popular misinformation websites compared to 2 years ago.
@@
To further support our finding, we explore how business relationships with misinformation websites have been affected based on their popularity. As a case study, we focus on Google, the most popular ad network (papadogiannakis2023pooling, ). We group misinformation websites into equal-sized buckets based on their network traffic, expressed in monthly visits during January 2024. As in previous figures, to increase our confidence on the validity of these business relationships, we cross-reference these accounts with those that Google reports in their sellers.json and only report those accounts that match. That is, we only study advertising accounts that Google agrees are valid within their ad system. We then compute the difference in the misinformation websites that work with these ad networks between December 2021 and January 2024. We express this difference as a percentage increase $D=\frac{A-B}{|B|}\cdot 100$, where $A$ (i.e., After) is the number of misinformation websites having a direct business relationship with the ad network on 2024/01, while $B$ (i.e., Before) is for 2022/01. We plot our findings in Figure 7. The major difference in the number of misinformation websites that no longer have a verified account with Google is observed for less popular websites (i.e., left-most bar of the plot). For example, approximately 96% of the misinformation websites that are ranked on the bottom 20% based on their network traffic no longer have a direct business relationship with Google. On the other hand, we observe that there is almost no difference for popular misinformation websites (i.e., right-most bar of the plot). Thus, we observe a discrepancy on how popular misinformation websites are treated, compared to unpopular ones.
@@
Inspired by the finding of Figure 7, we study if there is a correlation between the network traffic of a misinformation website and whether an ad network stopped working with this website. For this, we make use of the traffic labels that MBFC assigns to news websites. MBFC estimates traffic data through the SimilarWeb platform, but also takes into consideration the number of subscribers for print media and market size for TV or Radio entities (mbfcMethodology, ). The labels assigned to each website are “Minimal”, “Low”, “Medium” or “High”. We focus on the ad networks of Figure 3 that showed the greatest decrease in the number of misinformation websites they work with. We find that there is indeed a correlation between the network traffic of a misinformation website and if it declares a business contract with an ad network after two years. We present our findings in Table 1. The Spearman correlation coefficient is $\sim$0.6 with a $p$-value close to zero for all of the three ad networks we study. In fact, we observe that over 70% of misinformation websites with minimal network traffic no longer have a business relationship with these ad networks after two years.
Recent work has demonstrated that misinformation websites often use dark pooling in order to pool their ad inventory with that of lawful and acceptable websites (papadogiannakis2023pooling, ; vekaria2023inventory, ; vekaria2024turning, ). In an attempt to prove explicit relationships between popular misinformation websites and popular ad networks, we study their sellers.json files. In Figure 7, we plot the number of misinformation websites from our list of  2,500 misinformation websites that are explicitly stated inside sellers.json files of popular ad systems. This is a definitive proof that ad networks still fund misinformation websites because they explicitly name them as verified ad inventory sellers.
@@
Table 1. Correlation between misinformation websites traffic and progress of direct business contracts with ad networks. The last column indicates the percentage of websites with minimal traffic that no longer have a business relationships after 2 years.
Figure 9. Direct business contracts with misinformation websites on July 2024 based on ads.txt and sellers.json. Ad networks depicted in red are CoP signatories (copodSignatories, ).
We observe that the number of misinformation websites explicitly disclosed in the sellers.json files of the most prestigious and popular ad networks is somewhat stable. It is evident that the behavior of the top ad networks has not changed significantly over the last two years, even after the Code of Practice on Disinformation was signed. In fact, studying the standard deviation of the number of popular misinformation websites in sellers.json files (less than 2.5) shows that almost all networks have not stopped doing business with the misinformation websites.
To our surprise, we find that even today, there are a lot of ad networks working with certain misinformation websites that one can hardly argue they are legitimate. For example, we find that there are at least 10 of the most popular ad networks (including Google, Amazon, AppNexus and Pubmatic) that have an explicit and direct business relationship with worldlifestyle.com. According to MediaBias/FaceCheck, WordLifestyle is “a strong pseudoscience website based on the promotion of false or misleading claims regarding science and health” and it has a low factual reporting score “due to a lack of sourcing and blatant clickbait abuse to generate advertising revenue” (mbfcWorldLifeStyle, ). Even such websites that are known to take advantage of the ad ecosystem in order to monetize misinformation still have direct business contracts with popular ad networks. In fact, by manually visiting this website, we find that Google indeed serves ads to visitors.
In addition to this, we find that the website of “One America News Network” (OANN) is an approved ad inventory seller within the most popular ad networks (including Google and Appnexus, signatories of the Strengthened Code of Practice on Disinformation). OANN is considered a far-right biased news website that promotes propaganda, conspiracy theories and has over 10 failed fact checks (mbfcONNN, ). According to SimilarWeb, this website achieves over 2.5M monthly visits, and in conjunction with the fact that almost 95% of the visitors come from the United States, an advertising valuable country, it can generate substantial ad revenue.
Finding 1: Major ad networks still accept business relationships with misinformation websites which account for the majority of the misinformation traffic. After the CoP these ad networks have ceased their collaboration mainly with the unpopular misinformation web sites.
@@ 5. Financial Incentives
The analysis of Section 4 suggests that misinformation websites still rely on popular ad networks to monetize their content. Out of the 1,132 misinformation websites that served an ads.txt file on January 2024, we find that 1,126 of them (99.4%) disclose at least one direct business relationship with at least one ad network. The insightful reader may recognize that this information is based on ads.txt files reported by misinformation publishers themselves and assume that since these publishers generate fake content, they might also generate fake ads.txt files. To address this, we cross reference the accounts declared in ads.txt files with those found in sellers.json ad networks, just like in Section 4. We find that there exist 985 (87%) misinformation websites with at least one substantiated business relationship with an ad network. That is, for 985 misinformation websites there exists an ad network that confirms the publisher account exists within its network and that it indeed belongs to a publisher. This demonstrates that misinformation websites are still able to get approved by ad networks and monetize their content through ad revenue.
5.1. Direct Business Relationships
We set out to explore what is the state of the advertising ecosystem two years after the Strengthened CoP was signed. We collect the ads.txt files of misinformation websites and the sellers.json files of ad networks on July 2024. We plot in Figure 9 the number of direct business relationships among misinformation websites and ad networks. This relationships are extracted based on the ad accounts that both publishers and ad networks agree are valid through ads.txt and sellers.json files respectively. We observe that the majority of misinformation websites have a direct contract with only a handful of ad networks. However, there exist misinformation websites that form business contracts with tens or even hundreds of ad networks. One such example is the website whatreallyhappened.com. This website serves an ads.txt file of over 8500 lines that contains verified DIRECT entries for 196 distinct ad networks. In fact, the identifiers in this file are organized and labeled accordingly, indicating that this website participates in various (dark) pools (papadogiannakis2023pooling, ; vekaria2023inventory, ).
@@ 5.1. Direct Business Relationships
Next, we plot in Figure 9 the direct business relationships of top ad networks with misinformation websites. We observe that for most ad networks, the behavior is similar and have a DIRECT business relationship with about 16.9%-58.1% of misinformation websites whose ads.txt files we were able to obtain. Even though these values might seem small, the underlying implications are significant. They indicate that the most popular ad networks have a direct business contract with a large percentage of misinformation websites, and that they facilitate the proliferation of misinformation content through ad revenue. In fact, event the smallest value in Figure 9 suggests that ad networks have authorized at least 1 out of 5 misinformation websites that serve an ads.txt file. The two ad systems that stand out are Google, and Underdog Media that controls udmserve.net. These two ad networks have a direct business relationship with over 400 misinformation websites according to their sellers.json files. Interestingly, both ad networks disclose in their guidelines that they do not allow the monetization of misinformation content or content that makes false claims (googlePublisherGuidelines, ; underdogMediaPublisherGuidelines, ), but evidently they are not able to filter out all of these websites. Altogether, there are hundreds of misinformation websites that still have direct business relationships with legitimate and popular advertising networks and are able to generate revenue from dis-/mis-infomration content. The accounts used to sell the ad inventory of misinformation websites are recognized by popular ad networks.
@@ 5.2. News Websites Credibility
To further support our finding that ad networks still work with misinformation websites to a large extent, we utilize the Science Feedback dataset on news websites credibility (scienceFeedbackDataset, ). The researchers behind this dataset have accumulated a list of over 20K news websites from more than 100 countries. For each news website they assign a score reflecting its credibility, aggregating pre-existing and publicly available evaluations from trustworthy entities (i.e., academic researchers, fact-checking organizations, expert raters).
We access the original list and successfully fetch the ads.txt files of $\sim$7K news websites in July 2024. We perform a similar analysis as described earlier, but here, we make use of the list’s original classification system and focus on websites labeled as “Generally unreliable” and “Least reliable”. We are able to extract 2,459 unreliable news websites with at least one DIRECT record in their ads.txt files and present our findings in Figure 11. We discover a similar behavior with our previous findings. Popular and legitimate ad networks have a direct business contract with hundreds of unreliable news websites and this relationship is recognized by both the websites and ad networks.
Finally, in Figure 11 we focus on the signatories of the Code of Practice on Disinformation and explore their direct business relationships with news websites of various credibility scores. We find that these ad networks work with all news websites and do not seem to have a distinctively different policy for unreliable news websites. In the case of Google, we find that it works with almost as many of the least reliable websites (percentage wise) as with the most reliable.
@@ 6. Served Advertisements
In previous sections, we demonstrate that the most popular and legitimate ad networks still have business relationships with misinformation websites, and that the CoP did not have a substantial impact on the monetization of misinformation. To validate our observations, we study the ads that are actually served on misinformation websites. Our goal is to explore to what extent ad networks still monetize misinformation content and if reputable advertisers might end up next to misinformation content.
Inspired by previous findings that misinformation websites might have a short longevity (chalkiadakis2021rise, ), we aim to refine our list by accumulating only misinformation websites which are still active. Towards this extent, we perform a manual analysis using independent reviewers. We start from the extended list of 2,469 misinformation websites and using an automated crawler we take screenshots of the entire landing page of each website. Our crawler was able to successfully access 2,049 websites and capture screenshots. Next, two independent reviewers coming from a computer science background (not the authors of this work) evaluate which websites are still active based on the content of the landing pages. Reviewers are allowed classify each website as either “Active” or “Inactive”. Moreover, reviewers are allowed to manually visit websites and navigate through its pages if they are not certain which label to assign simply based on the screenshot of the landing page. The two reviewers are independent of each other and completed the task in a different environment, without influencing one another.
Both reviewers evaluated evaluated the entire list of 2,469 misinformation websites and agreed to include 1,889 websites, agreed to exclude 568 websites and disagreed on 11 websites. The reviewers agreed in 99.55% of websites and had a Cohen Kappa score of 0.99. Note that a high inter-rater agreement score is expected in this context because deciding if a website is active is an easy task for a human evaluator. We decide to only accept websites that both raters agree are active and end up with a list of 1,889 active misinformation websites. We make this list of active misinformation websites publicly available (openSourceData, ). The analysis of the following sections is based on this list of active websites only. We then analyze the ads found in misinformation websites, based on the methodology described in Section 3.3.

### v2

@@ 4. Regulation Impact
Figure 9. Direct business contracts with misinformation websites on July 2024 based on ads.txt and sellers.json. Red bars represent CoP signatories (Comission, 2022b).
We set out to explore if the Strengthened Code of Practice on Disinformation has had a meaningful impact on the monetization of misinformation content. We expect that the ecosystem will be improving because multiple signatories have committed to cutting the advertising revenue of disinformation spreaders (Comission, 2022a). Inspired by prior work (Papadogiannakis et al., 2023), we study how direct business relationships of misinformation websites with ad networks have evolved over time.
We utilize the dataset of ads.txt and sellers.json files described in Section 3.2, as well as the one collected in (Papadogiannakis et al., 2023) for news websites in December 2021. Arguably, publishers of fake news content could easily lie in their ads.txt files (Papadogiannakis et al., 2025; Vekaria et al., 2024a). To discover the actual business relationships, we focus on information coming from both the publishers and the ad systems. That is, we cross-reference the claims that publishers make in their ads.txt files with those of trustworthy ad networks in sellers.json files. We argue that since these are conflicting sides of the same coin, the intersection of their claims is the ground truth. For each distinct DIRECT ads.txt entry, we review whether the respective publisher ID is also disclosed by the respective ad network. For instance, if examplefakenews.com has the entry adnetwork.com, 12345, DIRECT in its ads.txt file, we confirm that adnetwork.com also discloses the identifier 12345 in its sellers.json file. Again, we focus only on ads.txt records labeled as DIRECT because they indicate a direct business contract between the misinformation publisher and the advertising system. First, we study the behavior of the top ad networks and how their business relationships with misinformation websites have evolved over a period of 2 years. We focus on the intersection of these datasets and find that there exist 461 misinformation websites with an ads.txt file in both datasets and with at least one business relationship also disclosed by the respective ad network. We plot our findings in Figure 3.
At first glance, most ad networks (as of January 2024) still work with misinformation websites, with only minor reductions. 6 out of the 9 most popular ad networks (i.e., rubiconproject.com to openx.com in Figure 3) have only blocked a very small percentage of misinformation websites (i.e., up to 5.4%). However, some ad systems, including Google, Lijit, and Amazon, have stopped working with a significant portion of misinformation sites, dropping 51%, 43%, and 37%, respectively, compared to two years ago. While the drop in ad network relationships with misinformation websites aligns with the CoP, these websites represent a small fraction of total misinformation traffic. To explore this, we use SimilarWeb (LTD, 2023) to gather traffic data for 95% of the misinformation websites for which we have collected their ads.txt files, before and after the CoP. In Figure 3, we plot the distribution of their network traffic (number of visits on desktop and mobile) in January 2024. The $x$-axis is in logarithmic scale, suggesting that there exist a lot of misinformation websites with low network traffic and a few highly popular that are able to attract millions of monthly visitors. We split these websites on the 80th percentile based on the observation that the top 20% are popular websites, while the rest are unpopular. [^3] The split 80-20 could have been made as 85-15, or 75-25. The point here is not to define an optimal threshold, but to create two categories: the “popular” misinformation websites and the “unpopular” ones. The 80th percentile corresponds to a network traffic of approximately 290K monthly visits. Consequently, we find 88 misinformation websites which we label as “popular”, and 351 labeled as “unpopular”.
We replicate the analysis from Figure 3 to track the evolution of ad network relationships with popular and unpopular misinformation sites in Figures 6 and 6, respectively. These figures paint a clear picture of the ad ecosystem. Even though it seems that some ad networks no longer work with misinformation websites, they do so only for the unpopular ones. Some of these websites have hardly any visitors and their effect on the dissemination of misinformation is limited, deeming them potentially insignificant. However, for high-traffic misinformation sites attracting millions of visits, ad networks continue to partner with them, even a year and half after the CoP was signed (Figure 6). Some top networks have severed ties with a small portion of the most popular misinformation websites, while others, like Amazon and Microsoft (AppNexus), now work with more popular misinformation websites than two years ago.
@@
To further support our finding, we explore how business relationships with misinformation websites have been affected based on their popularity. As a case study, we focus on Google, the most popular ad network (Papadogiannakis et al., 2025). We group misinformation websites into equal-sized buckets based on their monthly traffic in January 2024. To validate these business relationships, we cross-reference them with Google’s sellers.json accounts, only reporting those that match. We then compute the difference in misinformation websites working with Google between December 2021 and January 2024. We express this difference as a percentage decrease $D=\frac{B-A}{|B|}\cdot 100$, where $A$ (i.e., After) is the number of misinformation websites having a direct business relationship with Google on 2024/01, and $B$ (i.e., Before) is for 2022/01. Our findings, shown in Figure 6, reveal a significant drop in verified accounts for less popular websites (left-most bar of the plot), with 96% of websites ranked in the bottom 20% based on traffic, no longer having a contract with Google. On the other hand, we observe almost no difference for popular misinformation websites (i.e., right-most bar of the plot). Thus, we observe a discrepancy on how popular misinformation websites are treated, compared to unpopular ones.
@@
Building on Figure 6, we examine the correlation between a misinformation website’s network traffic and whether an ad network stops working with it. We make use of the traffic labels that MBFC assigns to news websites (“Minimal”, “Low”, “Medium” or “High”). MBFC estimates traffic data through the SimilarWeb platform, but also takes into consideration the number of subscribers for print media and market size for TV/Radio entities (Check, 2024b). We focus on the ad networks of Figure 3 that showed the greatest decrease in misinformation websites. We find a correlation between the network traffic of a misinformation website and if it declares a business contract with an ad network after two years. We present our findings in Table 1, with a Spearman correlation coefficient of $\sim$0.6 and a $p$-value close to zero. Over 70% of misinformation websites with minimal network traffic no longer have a business relationship with these ad networks after two years. We study how the popularity of misinformation websites has changed over time in Appendix B.
@@
Table 1. Correlation between misinformation websites traffic and progress of direct ad networks business contracts. The last column shows the percentage of low-traffic websites that no longer have business relationships after 2 years.
Recent work has demonstrated that misinformation websites often use dark pooling in order to pool their ad inventory with that of lawful and acceptable websites (Papadogiannakis et al., 2025; Vekaria et al., 2024a, b). In an attempt to prove explicit relationships between popular misinformation websites and popular ad networks, we study their sellers.json files. In Figure 9, we plot the number of misinformation websites from our list of $\sim$2,500 misinformation websites that are explicitly stated inside sellers.json files of popular ad systems. This is a definitive proof that ad networks still fund misinformation websites because they explicitly name them as verified ad inventory sellers.
We observe that the number of misinformation websites explicitly disclosed in the sellers.json files of the most prestigious and popular ad networks is somewhat stable. It is evident that the behavior of the top ad networks has not changed significantly over the last two years, even after the Code of Practice on Disinformation was signed. In fact, studying the standard deviation of the number of popular misinformation websites in sellers.json files (less than 2.5) shows that almost all networks have not stopped doing business with the misinformation websites. We provide examples of such relationships in Appendix C.
Finding 1: Major ad networks still accept business relationships with misinformation websites, which account for the majority of the misinformation traffic. After the CoP, these ad networks have ceased their collaboration mainly with the unpopular misinformation websites.
@@ 5. Financial Incentives
The analysis of Section 4 suggests that misinformation websites still rely on popular ad networks to monetize their content. We cross-reference the accounts declared in ads.txt files with those found in sellers.json ad networks, just like in Section 4 and find that there exist 985 (87%) misinformation websites with at least one substantiated business relationship with an ad network. That is, for 985 misinformation websites, there exists an ad network that confirms the publisher account exists within its network and that it indeed belongs to a publisher. This demonstrates that misinformation websites are still able to get approved by ad networks and monetize their content through ad revenue.
@@
5.1. Direct Business Relationships
We set out to explore what is the state of the advertising ecosystem two years after the Strengthened CoP was signed. We collect the ads.txt files of misinformation websites and the sellers.json files of ad networks on July 2024. We plot in Figure 9 the number of direct business relationships among misinformation websites and ad networks. These relationships are extracted based on the ad accounts that both publishers and ad networks agree are valid through ads.txt and sellers.json files, respectively. We observe that the majority of misinformation websites have a direct contract with only a handful of ad networks. However, there exist misinformation websites that form business contracts with tens or even hundreds of ad networks. One such example is the website whatreallyhappened.com. This website serves an ads.txt file of over 8500 lines that contains verified DIRECT entries for 196 distinct ad networks. In fact, the identifiers in this file are organized and labeled accordingly, indicating the presence of (dark) pools (Papadogiannakis et al., 2025; Vekaria et al., 2024a).
Next, we plot in Figure 9 the direct business relationships of top ad networks with misinformation websites. We observe that for most ad networks, the behavior is similar; they have a DIRECT business relationship with about 16.9%-58.1% of misinformation websites whose ads.txt files we were able to obtain. Even though these values might seem small, the underlying implications are significant. They indicate that the most popular ad networks have a direct business contract with a large percentage of misinformation websites, and that they facilitate the proliferation of misinformation content through ad revenue. In fact, event the smallest value in Figure 9 suggests that ad networks have authorized at least 1 out of 5 misinformation websites that serve an ads.txt file. The two ad systems that stand out are Google, and Underdog Media that controls udmserve.net. These ad networks directly partner with 400+ misinformation websites according to their sellers.json files. Even though their own guidelines prohibit monetizing false content (Google, 2024; Media, 2024), they fail to filter out all such websites. Altogether, there are hundreds of misinformation websites that still have direct business relationships with legitimate and popular ad networks and are able to generate revenue from dis-/mis-information content. The accounts used to sell the ad inventory of these websites are recognized by popular ad networks.
@@ 5.2. News Websites Credibility
To further support our finding that ad networks still work with misinformation websites to a large extent, we utilize the Science Feedback dataset on news websites credibility (Feedback, 2023). The researchers behind this dataset have accumulated a list of over 20K news websites from more than 100 countries. For each news website they assign a score reflecting its credibility, aggregating pre-existing and publicly available evaluations from trustworthy entities (i.e., academic researchers, fact-checking organizations, expert raters).
We access the original list and successfully fetch the ads.txt files of $\sim$7K news websites in July 2024. We perform a similar analysis as described earlier, but here, we make use of the list’s original classification system and focus on websites labeled as “Generally unreliable” and “Least reliable”. We are able to extract 2,459 unreliable news websites with at least one DIRECT record in their ads.txt files and present our findings in Figure 11. We discover a similar behavior with our previous findings. Popular and legitimate ad networks have a direct business contract with hundreds of unreliable news websites, and this relationship is recognized by both the websites and ad networks. Finally, in Figure 11 we focus on the signatories of the Code of Practice on Disinformation and explore their direct business relationships with news websites of various credibility scores. We find that these ad networks work with all news websites and do not seem to have a distinctively different policy for unreliable news websites. In the case of Google, we find that it works with almost as many of the least reliable websites (percentage wise) as with the most reliable.
@@ 6. Served Advertisements
In previous sections, we demonstrate that the most popular ad networks still have business relationships with misinformation websites, and that the CoP did not have a substantial impact on the monetization of misinformation. To validate our observations, we study ads that are actually served on misinformation websites. Our goal is to explore to what extent ad networks still monetize misinformation content and if reputable advertisers might end up next to misinformation content. This analysis follows the methodology described in Section 3.3. Since misinformation websites are short-lived (Chalkiadakis et al., 2021), we focus only on active ones, detailed in Appendix D.
6.1. Ad Networks
Using data coming from 6 distinct user profiles (2 genders and 3 countries) and one clean user profile, we detect ads from 2,480 advertisers in 392 misinformation websites. We attribute the large number of distinct advertisers to the use of different vantage points (i.e., 3 geographic locations), thus being served ads from local brands.
First, we study the ad networks that deliver ads to misinformation websites. We achieve this by studying the ad URLs the crawler detected. We analyze these ad URLs and extract the domain from them, grouping them on the eTLD+1 level. For example, ads served by adclick.g.doubleclick.net and googleads.g.doubleclick.net are all grouped under the doubleclick.net domain. To uncover the actual entities serving ads, we group domains based on the companies that own these domains. For example, ads served by doubleclick.net and googleadservices.com are all attributed to Google LLC. To identify entities that operate each domain, we utilize the DuckDuckGo Tracker Radar dataset (DuckDuckGo, 2024). We recognize that ad inventory (re)sales may involve intermediaries, but this is beyond the scope of this work. We summarize our findings in Table 2.